### v1

# Optimal control transport of neutral atoms in optical tweezers at finite temperature

Abstract
The transport of neutral atoms in Rydberg quantum computers is a crucial step of the initial arrangement of the grid as well as to the dynamic connectivity, recently successfully demonstrated [1]. We study the application of optimal control and the quantum speed limit for the transport of neutral atoms in optical tweezers at finite temperatures and analyze how laser noise affects transport fidelity. Open-loop optimal control significantly enhances transport fidelity, achieving an improvement up to 89% for the lowest analyzed temperature of $1\,\mu$K for a distance of three micrometers. Furthermore, we simulate how the transport fidelity behaves in release-and-capture measurements, which are realizable in the experiment to estimate transport efficiency and implement closed-loop optimal control.
Neutral atom quantum processing units based on optical tweezers [2, 3, 4, 5, 6, 7, 8] open the possibility to modify the connectivity of qubits at runtime of the algorithm. In the pioneering experiments in Ref. [9, 1], coherent transport of atoms through the movement of optical tweezers has been demonstrated. Ensuring high-fidelity atom movement is crucial not only for enabling non-local connectivity but also for the necessary reshuffling of atoms in the initial array arrangement, after the stochastic loading of atoms within the optical tweezers [10, 11, 12, 13, 14]. A fast high-fidelity transport can be achieved via quantum optimal control (QOC) [15, 16, 17, 18]. Optimal control has emerged as an essential tool for optimizing quantum gates as well [19, 20, 21, 1]. QOC allows for either open-loop optimization using numerical simulations or closed-loop optimization directly integrated with the experimental setup. A variety of optimal control methods is available via open-source package [22] as the dressed chopped random basis (dCRAB) algorithm [23, 24, 25].
[…]
Figure 1: Setup for atoms transport. (a) An array of neutral atoms after the initial stochastic loading. Empty traps are switched off and atom transport along one row of the array is possible via optical tweezer generated by acousto-optical deflectors (AODs). (b-c) We aim to move the atom at position $r_{0}=0\mu\mathrm{m}$ to the position $r_{\mathrm{f}}=3\mu\mathrm{m}$, a vacant spot in the example. (d) We perform open-loop optimal control via the time-averaged figure of merit $J_{\mathrm{avg}}$; closed-loop optimal control is anticipated by taking into account additional experimental constraints in our simulation, especially on the pulse shape $r(t)$.
We design an open-loop optimal transport of neutral atoms via optical tweezer at finite temperature to closely replicate experimental conditions: indeed, in typical experiments, atoms are trapped in Gaussian-shape potential formed by optical tweezers and each atom can be represented as a thermal state at a temperature approximately around $10\mu\mathrm{K}$ [27]. After the stochastic loading of the atoms, empty traps are switched off, as depicted in Fig. 1(a). We have control to move tweezers within a given row through acousto-optical deflectors (AODs). We reshuffle the atoms to occupy the vacant positions, thereby achieving the desired layout. The movement must be executed at sufficient fidelity and speed to prevent loss and avoid any temperature alterations [28]. We take into consideration a variety of experimental constraints for optical tweezers implemented via acousto-optical deflectors, e.g., a finite frequency bandwidth, and piecewise quadratic pulse shapes, both motivated by the experiment in Ref. [27, 29, 30]. We also introduce noise into the modeling of the tweezer trap, i.e. for its depth, waist, and position. We find that, with our setup, the quantum speed limit for moving a Strontium-88 atom over a distance of three micrometers is around $10\,\mu$s. Furthermore, at the lowest analyzed temperature, a notable enhancement in the figure of merit is observed, reaching a factor of 89%. We examine the impact of laser noise on the trap on the figure of merit. The release-and-capture protocol is the standard approach to verify the quality of the atom cooling [27], which also applies to the evaluation of the heating during transport. We predict which level of precision of the figure of merit can be observed experimentally by running the release-and-capture simulations after the optimal pulses.
We first describe the open-loop optimal control setup in Sec. I. Then, we turn to the analysis of the quantum speed limit in Sec. II, followed by simulations with noise and an explanation of the results in release-and-capture measurements based on these simulations. We conclude the results in Sec. III.
I Optimal control setup
We focus on the general problem of moving one atom from position $r_{0}$ to $r_{\mathrm{f}}$, e.g., from the configuration shown in Fig. 1(b) to that in Fig. 1(c). Each atom is in a thermal state
$\rho(T)$ $=\sum_{i=1}^{n_{\rm\scriptscriptstyle c}}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$ (1)
where $p_{i}=p_{i}(T)=\mathrm{e}^{-E_{i}/k_{b}T}$ are the Boltzmann weights at temperature $T$. We introduce a cutoff in the number of states $n_{\rm\scriptscriptstyle c}$ to represent the finite temperature density matrix. The convergence of the results with the cutoff $n_{\rm\scriptscriptstyle c}$ is discussed in appendix  A. We model the optical tweezer traps as Gaussian potentials of the following form:
$V(t)=U_{0}\mathrm{e}^{\frac{-2r(t)^{2}}{w_{0}^{2}}}\,,$ (2)
where $U_{0}$ is the trap depth, $w_{0}$ the beam waist and $r$ the time-dependent center of the trap [11, 6, 29]. In the experiment, the pulse $r(t)$ tuned to move the atom is generated via arbitrary waveform generators (AWG), which imposes constraints on the available pulse shapes, such as the piecewise quadratic shape [31]. Additionally, we consider laser noise effects and finite bandwidth during the optimization process. We describe the numerical simulation for computing the time-evolution of the thermal state under time-dependent pulses in appendix A and integrate quantum optimal control using the open-source suite QuOCS [22]. We employ the dCRAB algorithm [25, 24]: the key idea is to minimize a figure of merit, i.e., $J(r)$, by expanding the control pulse $r(t)$ as a sum of truncated randomized basis $r_{i}(t)$. In more detail, for a super-iteration $j$, the control pulse is constructed by finding the optimal set of coefficients $c_{i}$:
$r^{j}(t)=c_{0}^{j}r^{j-1}(t)+\sum_{i=1}^{N_{c}}c_{i}^{j}r_{i}^{j}(t)\,.$ (3)
A common selection for basis decomposition is the use of trigonometric functions, specifically the Fourier basis; however, the choice of bases can vary depending on the nature of the control problem, with certain bases proving more effective than others [32]. Specifically, we employed the sinc basis for optimizing the pulse $r(t)$. The figure of merit is the infidelity between the density matrices at the end of the pulse $\rho_{\mathrm{f}}(r)$ and the density matrix of the initial finite-temperature state shifted to the final position $\rho_{0}(r_{\mathrm{f}})$, i.e.,
[…]
In the following analysis, we target Strontium-88 atoms which have a mass of $m=1.46\times 10^{-25}\,\text{kg}$ and we fix a distance between the traps in the array of $d=3\mu\text{m}$. For the Gaussian-shaped trap, we assume a trap depth of $U_{0}=-1\text{mK}$ with a beam waist $w_{0}=0.5\mu\text{m}$. The coordinate in position-space is set via the frequency of the AOD, where the conversion is $13\mu\text{m}=$ 1MHz. The noise for the position, waist, and depth of the Gaussian potential is modeled as follows: the amplitude of the laser controlling the trap depth exhibits a noise $\delta S$ with a spectrum following a power-law distribution given by $S(f)=A_{L}/f^{1/2}$, i.e., the relative intensity noise (RIN). We know from measurement [33] that the amplitude is about $A_{L}=10^{-11}$. The noise of the laser itself is not significant for the transport, but the beam is typically modified by fiber couplings, the optical system, and especially the AOD; the AOD with its raise time of $6\mu$s modifies the beam. This is considered as additional noise $\delta U_{0}$ in the trap depth, introduced by random sinusoidal fluctuations with an amplitude of 1% and random frequencies below 100100100\,100kHz, as well as random sinusoidal fluctuations with an amplitude of 5% and random frequencies exceeding 100100100\,100kHz. For the trap waist, we assume sinusoidal noise $\delta\omega_{0}$ with a frequency of $1/6\,\mu$s motivated by the raise time of the AOD and an amplitude of 1%. Regarding the trap position, we assume sinusoidal noise $\delta r$ with an amplitude of $0.01\,\mu$m and random frequencies ranging between $50-150\,$MHz.
II Time-optimal pulse for atom transport
First, we assess the performance of a pulse $r_{\mathrm{pq}}(t)$ matching the AWG’s constraint, specifically a piecewise quadratic pulse of the form:
$r_{\mathrm{pq}}(t)=\left\{\begin{aligned} &r_{0}+2\frac{r_{\mathrm{f}}-r_{0}}{t_{\mathrm{p}}^{2}}t^{2}&\quad&t\leq t_{\mathrm{p}}/2\\ &r_{\mathrm{f}}-2\frac{r_{\mathrm{f}}-r_{0}}{t_{\mathrm{p}}^{2}}(t-t_{\mathrm{p}})^{2}&\quad&t>t_{\mathrm{p}}/2\end{aligned}\right.$ (5)
where $t_{\mathrm{p}}$ is the duration of the pulse. Since the figure of merit defined in Eq. (4) oscillates at the end of the transport, we evolve the system for an additional fixed $t_{\mathrm{c}}=10\,\mu\text{s}$ to compute the time-average $J_{\mathrm{avg}}$. Thus, the evolution time is $t_{\mathrm{t}}=t_{\mathrm{p}}+t_{\mathrm{c}}$. The optimization workflow is summarized in Fig. 1(d), with an example of a piecewise quadratic pulse represented by the black line. We examine the time-average of the figure of merit, $J_{\mathrm{avg}}$, as a function of the pulse duration $t_{\mathrm{p}}$ for the piecewise quadratic pulse, considering various temperatures for the thermal state in Eq. (1), specifically for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K. The results are shown in Fig. 2(a). The eigenstates of the Gaussian potential lead to typical frequencies and timescales in the system, which is shown as periodic minima in the figure of merit; this result agrees with the observation for transport in optical lattices, see Ref. [34]. For the strontium-88 atom and for the selected trap depth, the periodicity is given by $2\tau\approx 10\,\mu$s with
$\tau=\frac{2\pi}{(E_{1}-E_{0})},$ (6)
where $E_{1}$ and $E_{0}$ are the eigenenergies of the first excited and ground state of the Gaussian trap, respectively. Using the piecewise quadratic pulse, we achieve a figure of merit $J_{\mathrm{avg}}<10^{-2}$ for $t_{\mathrm{p}}>20\,\mu$s, for the considered temperatures as shown in Fig. 2(a).
[…]
Figure 2: Figure of merit as a function of transport time for different temperatures. (a) For a fixed pulse duration $t_{\mathrm{p}}$, the center of the Gaussian-shaped trap is moved via a piecewise quadratic pulse $r_{{\mathrm{pq}}}$ for the distance of $3\mu$m. We observe periodic minima with a frequency of $2\tau=10\,\mu$s. When these minima occur, the pulse length $t_{\mathrm{p}}$ is precisely tuned to enable the wavefunction to complete its oscillation period. (b) The pulse $r_{{\mathrm{oc}}}$ is obtained with optimal control with $r_{{\mathrm{pq}}}$ as initial guess. The optimal control removes the dependency on the wavefunction oscillations and reaches a figure of merit $J_{\mathrm{avg}}$ which monotonically decreases with the pulse time $t_{p}$. We observe the quantum speed limit around $2\tau$. The figure of merit depends on the temperature $T$, i.e., higher temperatures result in a worse $J_{\mathrm{avg}}$ for the optimal pulse.
In Fig. 2(b) we now optimize the pulse shape via QOC: we analyze $J_{\mathrm{avg}}$ as a function of the optimal pulse with duration $t_{\mathrm{p}}$. The initial guess for the optimization is the piecewise quadratic pulse $r_{\mathrm{pq}}$. This data shows the typical feature of the quantum speed limit in an optimal control problem [35, 34]: we extract the quantum speed limit from Fig. 2(b) as $2\tau$ where the figure of merit does not improve anymore. This estimate is in good agreement with the timescale of the gap in Eq. (6). The optimal control optimization eliminates the figure of merit oscillations that were present in the piecewise quadratic case and results in $J_{\mathrm{avg}}=10^{-2}$ at around $t_{\mathrm{p}}=10-12\,\mu$s depending on the temperature $T$. We observe that the temperature significantly influences the figure of merit for atom transport. Higher temperatures lead to a worse figure of merit for the optimal pulse, with no room for further optimization. The relative improvement of the optimal control solution for the piecewise quadratic pulse is illustrated in Fig. 3 as a function of the pulse duration. The improvement holds particular significance at pulse duration $t_{\mathrm{p}}$ where minima were evident for $r_{\mathrm{pq}}$ and as close as possible to the quantum speed limit. The optimal control can explore interesting solutions in such scenarios [15]. Specifically at $t_{\mathrm{p}}=11\,\mu$s, the figure of merit can be improved by a factor of $89\%,58\%,66\%$ for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K, respectively. Furthermore, we verify that a comparable improvement is observed when we impose the experimental constraint on the pulse shape fitting it with piecewise quadratic segments, each spanning $10\,\mu$s.
[…]
Figure 3: Relative improvement of optimal control solution compared to piecewise quadratic pulse. We observe a substantial improvement with optimal control. This enhancement is particularly significant for pulses $r_{\mathrm{oc}}$ with a duration $t_{\mathrm{p}}$ aligning with minima observed in the piecewise quadratic solution $r_{\mathrm{pq}}$. Within the minima, we are interested in the region around the quantum speed limit. For example, we observe a significant advantage at $t_{\mathrm{p}}=11\,\mu$s, where the figure of merit improves by $89\%,58\%,66\%$ for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K, respectively.
[…]
Figure 4: Transport of neutral atoms under noisy optical tweezer. (a) The amplitude of the laser controlling the trap depth follows a noise $\delta S$ with spectrum following a power law $S(f)=A_{L}/f^{1/2}$ with $A_{L}=10^{-11}$. The measured noise of the laser itself is not strong enough to affect the transport, but the beam is modified by the optical fibers, the AOD, and the optical elements. Thus, we additionally consider a sinusoidal noise $\delta U_{0}$ for the trap depth. Moreover, we consider sinusoidal noise denoted by $\delta\omega_{0}$ for the trap waist and $\delta r$ for the position, respectively. (b) Time-dependent pulses for trap depth $U_{0}$ and waist $\omega_{0}$ are shown in the presence of noise. The sinusoidal noise $\delta r$ for the position $r$ is depicted in the bottom plot.
We then incorporate laser noise into our numerical simulation according to the description of Sec. I. We summarize the setup in Fig. 4(a) and show example pulses in panel (b). We calculate the figure of merit for the piecewise quadratic pulse $r_{\mathrm{pq}}$ with a fixed pulse duration of $t_{\mathrm{p}}=21\,\mu$s at $T=1\,\mu$K. Remember that this point corresponds to a minimum with $J_{\mathrm{avg}}=10^{-3}$ for $r_{\mathrm{pq}}$, as depicted in Fig. 2(a). Upon the introduction of noise, the infidelity is $J_{\mathrm{avg}}^{\mathrm{noise\,(pq)}}=0.015\pm 0.018$. This result is obtained by averaging over 100 simulations with different random noise instances. The substantial standard deviation suggests that outcomes heavily rely on the specific noise instance. Consequently, scenarios may vary, ranging from instances where noise is insufficient to affect results to those where the infidelity increases by one order of magnitude. We perform optimal control under noise and can improve the infidelity by 20%, i.e., $J_{\mathrm{avg}}^{\mathrm{noise\,(oc)}}=0.012\pm 0.013$.
In the experiment, release-and-capture measurements offer a method to estimate the temperature of the trapped atom [27]. This provides an opportunity to integrate closed-loop optimal control into the experiment by using the measurement from release-and-capture as a figure of merit for the optimal control instead of $J_{\mathrm{avg}}$. To test the sensitivity of the release-and-capture as a potential figure of merit, we simulate a release-and-capture measurements after applying the pulse $r(t)$. Here, we examine our final density matrices obtained after applying the piecewise quadratic pulse $r_{\mathrm{pq}}(t)$ or the optimal pulse $r_{\mathrm{oc}}(t)$. The sequence consists of turning off the trapping potential, allowing the wavefunction to freely evolve, turning on the potential again after a time $\tau_{\mathrm{rc}}$, and measuring the probability of having the atom still in the optical tweezer. While in the experiment there is a clear path towards detecting the remaining population of atoms via fluorescence, we have to define a corresponding figure of merit for the numerical simulations. Therefore, we calculate the probability of the wavefunction being inside the trap as the accumulated probability to be within the 35th and 65th percentiles of the integrated Gaussian potential. To amplify the information in the release-and-capture measurement, we move the atom $N_{t}$ times back and forth in the simulation and then measure; the backward movement is achieved by reversing the forward pulse. For example, the atom is moved ten times back and forth between the positions of $0\mu$m and $3\mu$m; with the last movement, it reaches again its final position of $3\mu$m with $N_{t}=21$. Looking at Fig. 5 with $N_{t}=1$, we observe that the measurements after applying the pulses $r_{\mathrm{pq}}(t)$ and $r_{\mathrm{oc}}(t)$ are indistinguishable; the release-and-capture is not as sensitive to errors as the numerical approach. The center panel of Fig. 5 shows that with $N_{t}=21$ we are now able to resolve the optimal control solution from the piecewise quadratic pulse. Going to $N_{t}=41$, the trend is preserved. Therefore, we propose for a closed-loop optimal control to apply the guess pulse multiple times before running release-and-capture measurements.
[…]
Figure 5: Release and capture postpulse measurements. For a temperature of $T=1\,\mu$K in the thermal state, we conduct release-and-capture measurements, i.e., the final density matrix, obtained after applying the piecewise quadratic pulse $N_{t}$ times, is evolved for a time $\tau_{\mathrm{rc}}$ with the trap turned off. The probability is computed as the accumulated probability for the wavefunction to be inside the 35th and 65th percentiles of the integrated Gaussian potential. We focus on the pulse with $t_{\mathrm{p}}=11\,\mu$s and compare the measurements for the piecewise quadratic pulse with the one optimized via optimal control. We do not observe any substantial difference between the measurements when we transport the atom only one time, i.e., for $N_{t}=1$. This absence of noticeable differences suggests that release-and-capture measurements are not sensitive to our figure of merit. However, when we transport the state many times, e.g., with $N_{t}=41$, we can observe a difference.
III Conclusion
We applied open-loop optimal control to transport atoms trapped in optical tweezers over a distance of one lattice site. The wave function is modeled as a thermal state, and an analysis is conducted for various temperatures. While we chose the parameters for strontium-88 atoms, the entire analysis is adaptable for other species of atoms. The optimal control is implemented using the open-source library QuOCS [22], and the dCRAB algorithm is utilized for pulse optimization [25]. We quantified the quality of the transport using the time-averaged overlap between density matrices. First, we simulated the transport with piecewise quadratic pulses, which then served as a benchmark for optimal control. For the piecewise quadratic pulse, we observed the expected behavior, i.e., local minima of the transport infidelity matching in their periodicity the eigenfrequencies of the Gaussian potential. Subsequently, we executed the optimal control optimization and achieved improvements of up to 89% for the lowest chosen temperature of $1\,\mu$K. We confirm that the results do not significantly change when we impose the constraint of having only piecewise quadratic pulse shapes in the optimal control. When we include noise on the trap depth, waist, and position for one data point, the infidelity increases on average by two orders of magnitude. With optimal control, we can get an improvement of 10%. Finally, we conducted release-and-capture measurements, potentially suitable for implementation in the experiment for closed-loop optimal control optimization. These measurements are sensitive to our figure of merit if the transport is performed several times.
In the next steps, the assumptions on the noise can be further refined with experimental data toward an even more refined open-loop optimal control. The closed-loop optimal control also has now an input in terms of speed-limit potential expectation of how to use release-and-capture measurements as a figure of merit.
Code and data availability -- The code to perform the optimal control is available as QuOCS [22]. The code for transport is available upon reasonable request.
Acknowledgments -- We thank our partners from the QRydDemo project Hans Peter Büchler, Florian Meinert, Tilman Pfau, and Jürgen Stuhler for useful discussions. We thank Aaron Götzelmann, Christian Hölzl, and Christoph Tresp for providing experimental data and discussing the experimental data and setup. We thank Atiye Abedinnia for the discussions. This project has received funding from the German Federal Ministry of Education and Research (BMBF) under the funding program quantum technologies -- from basic research to market -- with the grant QRydDemo. We acknowledge financial support from the Italian Ministry of University and Research (MUR) via PRIN2022 project TANQU, and the Department of Excellence grant 2023-2027 Quantum Frontiers; from the European Union via H2020 projects EuRyQa, TEXTAROSSA, and the Quantum Flagship project Pasquans2, the EU-QuantERA projects QuantHEP and T-NISQ, and from the World Class Research Infrastructure -- Quantum Computing and Simulation Center (QCSC) of Padova University, and the Italian National Centre on HPC, Big Data and Quantum Computing. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 40/575-1 FUGG (JUSTUS 2 cluster).
Appendix A Simulation approach for transport
The numerical simulation for the atom transport uses a split-step method similar to the proposal in Ref. [18]. The main idea of the method is to consider the Hamiltonian of the system as a sum of the potential and kinetic parts, i.e.,
$H$ $=K+V(x,t)\,,\qquad K=-\frac{h^{2}}{2m}\frac{\partial^{2}}{\partial x^{2}}\,.$ (7)
The time-dependent Schrödinger equation $\mathrm{i}h\frac{\partial}{\partial t}\ket{\psi}=(H+V)\ket{\psi}$ can then be approximated via a Trotter decomposition of $K$ and $V$ for the unitary propagator of one time-step $dt$ as
$\exp\left(\frac{-\mathrm{i}dt}{h}(K+V)\right)$
$\approx\exp\left(\frac{-\mathrm{i}dt}{h}K\right)\exp\left(\frac{-\mathrm{i}dt}{h}V\right)\,.$ (8)
For the density matrix $\rho$, we evolve each state |$\ket{\psi_{i}}$ from Eq. (1) independently and recombine them with their weights $p_{i}$. The error in the approximation depends on the commutator $[K,V]$ and scales with the timestep as $\mathcal{O}(dt^{2})$; the error for the complete time evolution scales as $\mathcal{O}(dt)$ for the first order Trotter decomposition. The key mechanism are Fourier transformations implemented via the FFT library to have both components $K$ and $V$ in their diagonal eigenbasis. The exponential, now a diagonal matrix, becomes cheap and the FFT carries the major part of the workload [36]. The spatial grid is static over the evolution; no moving grid is required for the distances considered here.
The convergence study targets the main parameters of the simulations, which are the discretization in time, in space, the range in space, and the cutoff for the finite temperature states. For each parameter, we specify the base size of a simulation as $n$; then, we compare the convergence with respect to simulations of a problem size of $2n$, $4n$, and $8n$, where the latter serves as a reference point. For the convergence in discretization in time and space, we take the piecewise quadratic scheme for a distance of $3\,\mu$m as a simulation in case of the ground state and the infidelity is calculated based on the overlap of the wave function
$\mathcal{F}(n)$ $=\left|\innerproduct{\psi(8\cdot n)}{\psi(s\cdot n)}\right|^{2},\,s\in\{1,2,4\}\,.$ (9)
The convergence study for the extent of the grid uses the same Eq. (9), but running a release-and-capture time evolution instead of a piecewise-quadratic transport of the atoms; the release-and-capture is more vulnerable to the extent of the system as there is no trapping potential during the evolution. For the cutoff in the excited states, we use the relation
$\mathcal{F}(n)$ $=\mathcal{F}(U\rho(s\cdot n)U^{\dagger},U\sigma(8\cdot n)U^{\dagger})$ (10)
$=\left(\sum_{i=1}^{s\cdot n}\sqrt{p_{i}q_{i}}\right)^{2}\,,$
where $p_{i}$ and $q_{i}$ are the probabilities of the eigenstates of $\rho$ and $\sigma$, respectively. Herein, we assume that we execute the same dynamics $U$, i.e., scenarios of optimal control finding two pulses $U$ and $V$ for different cutoffs are beyond this analysis and beyond our interest.
[…]
Figure 6: Convergence plots for various parameters. For every parameter, we show the convergence by comparing simulations at size $n$, $2n$, and $4n$ where $n$ is the default value of the parameter. The discretization of the spatial grid $dx$ shows a scaling of $\mathcal{O}(dx^{3})$, the one of the time grid scales as $\mathcal{O}(dt^{4})$. For the range, we increase the limits in the space $x_{\mathrm{max}}-x_{\mathrm{min}}$ at constant discretization $dx$. The cutoff $n_{c}$ is shown for $T=10\,\mu$K; we do not expect a linear scaling in log-log as the exponential distribution enters via the thermal state.
Table 1: Settings for convergence study. The default settings in terms of the size $n$ for simulation properties to be converged.
Property Default size $n$
Time discretization $dt$ $0.5\,\mu$s
Space discretization $dx$ $0.005\,\mu$m
Grid extension $x_{\mathrm{max}}-x_{\mathrm{min}}$ $5\,\mu$m
State cutoff $n_{c}$ 2
For the convergence study, we take the setup from Tab. 1 to define the default size $n$ of the property to be converged. The specific simulations for each parameter lead to the convergence according to Fig. 6 and the infidelity $\mathcal{I}$ is defined as $\mathcal{I}=1-\mathcal{F}$. The cutoff $n_{c}$ is shown for $T=10\,\mu$K. We observe convergence with $dx$ following a scaling of $\mathcal{O}(dx^{3})$. For the time discretization the scaling is $\mathcal{O}(dt^{4})$. Neither the simulations concerning the grid extension nor in the cutoff show a polynomial convergence. For the latter, the behavior was expected as the exponential of the thermal state goes into the convergence. According to the results of the convergence study, we adopt the following parameters for the simulation: a time discretization of $dt=0.1\,\mu$s, a spatial discretization of $dx=0.002\,\mu$m, a grid extension of $x_{\max}-x_{\min}=10\,\mu$m, and a state cutoff of $n_{c}=8$.